# Forum for Information Retrieval Evaluation, December 12-15, 2024, India

### v1

@@ Abstract
Language Identification (LI) is crucial for various natural language processing tasks, serving as a foundational step in applications such as sentiment analysis, machine translation, and information retrieval. In multilingual societies like India, particularly among the youth engaging on social media, text often exhibits code-mixing, blending local languages with English at different linguistic levels. This phenomenon presents formidable challenges for LI systems, especially when languages intermingle within single words. Dravidian languages, prevalent in southern India, possess rich morphological structures yet suffer from under-representation in digital platforms, leading to the adoption of Roman or hybrid scripts for communication. This paper introduces a prompt based method for a shared task aimed at addressing word-level LI challenges in Dravidian languages. In this work, we leveraged GPT-3.5 Turbo to understand whether the large language models is able to correctly classify words into correct categories. Our findings show that the Kannada model consistently outperformed the Tamil model across most metrics, indicating a higher accuracy and reliability in identifying and categorizing Kannada language instances. In contrast, the Tamil model showed moderate performance, particularly needing improvement in precision and recall.
@@ 1 Introduction
In the context of India, a country renowned for its linguistic diversity [2, 3, 4, 5], social media platforms reflect a vibrant mix of languages. Among the youth, in particular, there is a prevalent use of code-mixed text that blends local languages from the Dravidian language family with English. Dravidian languages, spoken predominantly in southern India, including languages like Kannada, Tamil, Malayalam, and Tulu, are characterized by rich morphological structures and diverse linguistic features. However, despite their significance, these languages face technological challenges, such as inadequate digital representation and script variations, which complicate language processing tasks like LI.
@@
In this work, we leveraged GPT-3.5 Turbo [6, 7, 8, 9, 10] to understand whether the large language models is able to correctly classify words into correct categories. We experiment with GPT at different temperature values namely 0.7, 0.8, and 0.9.
@@
We observe that for Tamil and Kannada, GPT models have significant room for improvement.
@@ 2 Related Work
Language Identification (LI) [6, 11, 12, 13, 14, 15] has been a crucial area of research within Natural Language Processing (NLP) due to its foundational role in various applications such as sentiment analysis, machine translation, and information retrieval. Traditional LI approaches [16, 17, 18, 19, 20, 21, 22] have primarily focused on monolingual or bilingual contexts, where clear boundaries between languages are assumed. However, these methods often struggle in multilingual and code-mixed environments, especially in regions like India, where linguistic diversity [23, 24, 25, 26, 27, 28, 29, 30] is high and social media usage reflects complex language practices.
Code-mixing [31, 32, 33, 34, 35, 36] presents unique challenges for LI systems. Early research in code-mixing focused on language pairs like English-Spanish or Hindi-English, where code-mixed texts predominantly used Roman scripts. Notable works explored the linguistic features of code-switched texts and highlighted the difficulties in segmenting and identifying languages at the word level. Similarly, Hindi-English code-mixed social media text, emphasizes the necessity for specialized LI models capable of handling intra-word language switches.
Dravidian languages [37, 38, 39, 40, 41] have been relatively underexplored in the context of LI, primarily due to the scarcity of annotated datasets and the complex morphological characteristics inherent to these languages. Previous efforts have developed initial datasets and models for LI in Dravidian languages; however, these models often fall short in handling code-mixed text, where Roman or hybrid scripts are employed. The Dravidian-CodeMix shared task aimed to address some of these gaps by introducing datasets for Tamil, Malayalam, and Kannada, which included code-mixed instances. Yet, the performance of models on these datasets indicated significant room for improvement, particularly in distinguishing between closely related languages and dialects.
Large Language Models (LLMs) [42, 43, 44, 45, 46, 47, 48, 49] like GPT-3 have shown promise in various NLP tasks, including LI. Previous works have demonstrated the capability of GPT-3 in performing zero-shot and few-shot learning, making it a potentially powerful tool for LI in resource-constrained settings. However, the application of LLMs [50, 51, 52, 53, 54, 55, 56, 57] to code-mixed and morphologically rich languages remains underexplored. Recent studies, have started to explore the use of transformers and pre-trained models for multilingual LI, but the effectiveness of these models in code-mixed Dravidian languages, particularly at the word level, requires further investigation.
@@ 3 Dataset
This shared task consists of four distinct datasets [58, 59, 60, 61, 62],:
@@
    Tulu Dataset: The Tulu dataset consists of 7,171 code-mixed sentences sourced from YouTube videos. These sentences have undergone preprocessing to eliminate non-textual characters and have been transliterated into Roman script. The dataset comprises 36,002 words categorized into six classes: ’Tulu’, ’Kannada’, ’English’, ’Mixed-language’, ’Name’, and ’Location’. The presence of mixed-language words poses significant challenges due to their dynamic and context-specific nature.
@@
    Kannada Dataset: This dataset includes 14,847 tokens represented in Roman script. It is classified into six categories: ’Kannada’, ’English’, ’Mixed-language’, ’Name’, ’Location’, and ’Other’. The dataset aims to enhance methods for language identification and categorization specifically tailored for Kannada-English code-mixed texts.
@@
    Tamil Dataset: The Tamil dataset consists of 17,568 tokens following a methodology similar to that of the Tulu and Kannada datasets. It is categorized into six classes and supports various natural language processing tasks specific to the Tamil language domain.
@@
    Malayalam Dataset: The Malayalam dataset comprises 25,035 tokens categorized into seven classes: ’Malayalam’, ’English’, ’Mixed’, ’Name’, ’Number’, ’Location’, and ’sym’ (for sentence boundaries). This dataset offers extensive coverage for natural language processing tasks and includes additional categories such as ’Number’ for numerical values, similar to the other datasets provided.
We participated in shared tasks based on two languages, namely, Kannada and Tamil. The test dataset size for Kannada is 2502. The test dataset size for Tamil is 2024.
@@ 5.1 Why Prompting?
Prompting [63] to solve a word-level classification problem often arises from the need to accurately identify and categorize individual words within texts that exhibit code-mixing or multilingual content. Here are some reasons why this problem is prompted:
@@
  • Accuracy in Language Processing: For effective natural language processing (NLP), identifying the language of each word enhances the accuracy of subsequent tasks [65, 66]. It ensures that language-specific models or algorithms are applied correctly.
@@ 5.2 Prompt Engineering-Based Approach
We used the GPT-3.5 Turbo (Figure 1) model via prompting [^1] https://platform.openai.com/docs/models/gpt-3-5-turbo to solve the classification task. We used GPT-3.5 Turbo in Zero-Shot mode via prompting. After the prompt is provided to the LLM, the following steps occur internally while generating the output. We list these steps, summarizing the prompting approach using GPT-3.5 Turbo:
@@ 6 Results
Metric Tamil Rank (Tamil) Kannada Rank (Kannada)
Macro F1 0.3312 10 0.4493 10
Macro Precision 0.3259 10 0.5474 10
Macro Recall 0.3657 10 0.4241 10
Weighted F1 0.7022 10 0.6725 10
Weighted Precision 0.7559 10 0.7191 10
Weighted Recall 0.6689 10 0.6994 10
Accuracy 0.6689 10 0.6994 10
Table 1: Comparison of various metrics for Word level identification in code-mixed languages in two languages-Tamil, Kannada. The team name is TextTitans and the username is roydanik18. The rank corresponding to a score is also provided.
Table 1 presents metrics comparing the performance of two language identification models, one for Tamil and the other for Kannada. Here’s a detailed discussion of each metric. For Tamil language, the macro F1 score is 0.3312. This suggests that the model achieves a balanced performance in terms of precision and recall for Tamil language identification. However, it indicates there is room for improvement in correctly identifying both positive and negative instances. For Kannada, the macro F1 score is 0.4493. This score is higher compared to Tamil, indicating a better overall balance between precision and recall for Kannada language identification. The model for Kannada performs better in correctly classifying instances across the dataset.
@@ 7 Conclusion
Moving forward, further refinements in model training and dataset augmentation could enhance the performance of language identification systems for both Tamil and Kannada. Future research efforts may focus on incorporating additional linguistic features, optimizing model architectures, and expanding datasets to include more diverse linguistic variations and challenges. In conclusion, this study underscores the importance of tailored approaches in language identification, particularly in multilingual settings like India where linguistic diversity is prominent. By advancing the capabilities of language identification models through innovative methodologies such as GPT-3.5 Turbo via prompting, we contribute to the broader goal of improving language processing technologies for diverse and under-resourced languages, fostering more accurate and inclusive natural language understanding systems.

### v2

@@ Abstract
Language Identification (LI) is crucial for various natural language processing tasks, serving as a foundational step in applications such as sentiment analysis, machine translation, and information retrieval. In multilingual societies like India, particularly among the youth engaging on social media, text often exhibits code-mixing, blending local languages with English at different linguistic levels. This phenomenon presents formidable challenges for LI systems, especially when languages intermingle within single words. Dravidian languages, prevalent in southern India, possess rich morphological structures yet suffer from under-representation in digital platforms, leading to the adoption of Roman or hybrid scripts for communication. This paper introduces a prompt based method for a shared task aimed at addressing word-level LI challenges in Dravidian languages. In this work, we leveraged GPT-3.5 Turbo to understand whether the large language models are able to classify words into correct categories correctly. Our findings show that the results on the Kannada dataset consistently outperformed the Tamil dataset across most metrics, indicating a higher accuracy and reliability in identifying and categorizing Kannada language instances. In contrast, the results on the Tamil dataset showed moderate performance, particularly needing improvement across all metrics.
@@ 1 Introduction
In the context of India, a country renowned for its linguistic diversity [2], social media platforms reflect a vibrant mix of languages. Among the youth, in particular, there is a prevalent use of code-mixed text that blends local languages from the Dravidian language family with English. Dravidian languages, spoken predominantly in southern India, including languages like Kannada, Tamil, Malayalam, and Tulu, are characterized by rich morphological structures and diverse linguistic features. However, despite their significance, these languages face technological challenges, such as inadequate digital representation and script variations, which complicate language processing tasks like LI.
@@
An example of the dataset structure and word categories (adapted from-https://sites.google.com/view/coli-dravidian-2024/datasets?authuser=0) for the task is shown in Figure  1.
To the best of our knowledge, there is no work which explores unsupervised approaches for language identification. In this work, we leveraged GPT-3.5 Turbo [3] to understand whether the large language models is able to correctly classify words into correct categories. We experiment with GPT at different temperature values namely 0.7, 0.8, and 0.9.
@@
Based on our experiments we observe that for Tamil and Kannada, GPT models have significant room for improvement.
Figure 1: Dataset structure and word categories for the task
@@ 2 Related Work
Language Identification (LI) [4, 5, 6] has been a crucial area of research within Natural Language Processing (NLP) due to its foundational role in various applications such as sentiment analysis, machine translation, and information retrieval. Traditional LI approaches [7, 8, 9, 10, 11] have primarily focused on monolingual or bilingual sentences, where clear boundaries between languages are assumed. However, these methods often struggle in multilingual and code-mixed environments, especially in regions like India, where linguistic diversity [12, 13, 14, 15, 16, 17] is high and social media usage reflects complex language practices.
Code-mixing [18, 19] presents unique challenges for LI systems. Early research in code-mixing focused on language pairs like English-Spanish or Hindi-English, where code-mixed texts predominantly used Roman scripts. Notable works explored the linguistic features of code-switched texts and highlighted the difficulties in segmenting and identifying languages at the word level. Similarly, Hindi-English code-mixed social media text, emphasizes the necessity for specialized LI models capable of handling intra-word language switches.
Dravidian languages [20] have been relatively underexplored in the context of LI, primarily due to the scarcity of annotated datasets and the complex morphological characteristics inherent to these languages. Previous efforts have developed initial datasets and models for LI in Dravidian languages; however, these models often fall short in handling code-mixed text, where Roman or hybrid scripts are employed. The Dravidian-CodeMix shared task aimed to address some of these gaps by introducing datasets for Tamil, Malayalam, and Kannada, which included code-mixed instances. Yet, the performance of models on these datasets indicated significant room for improvement, particularly in distinguishing between closely related languages and dialects.
Large Language Models (LLMs) [21, 22, 23, 24] like GPT-3 have shown promise in various NLP tasks, including LI. Previous works have demonstrated the capability of GPT-3 in performing zero-shot and few-shot learning, making it a potentially powerful tool for LI in resource-constrained settings. However, the application of LLMs [25, 26, 27, 28] to code-mixed and morphologically rich languages remains underexplored. Recent studies, have started to explore the use of transformers and pre-trained models for multilingual LI, but the effectiveness of these models in code-mixed Dravidian languages, particularly at the word level, requires further investigation.
@@ 3 Dataset
This shared task (adapted from  https://sites.google.com/view/coli-dravidian-2024/datasets?authuser=0) consists of four distinct datasets [29, 30, 31, 32, 33, 34, 35, 36, 33, 37]:
@@
    Tulu Dataset: This dataset is composed of 7,171 code-mixed sentences gathered from YouTube videos. These sentences have been cleaned to remove non-textual elements and transliterated into Roman script. The dataset contains a total of 36,002 words, which are organized into six categories: ’English’, ’Kannada’, ’Tulu’, ’Location’, ’Name’, and ’Mixed-language’. The dynamic and context-specific nature of mixed-language words presents notable challenges for processing.
@@
    Kannada Dataset: This Kannada dataset contains 14,847 tokens in Roman script and is divided into six categories: ’English’, ’Kannada’, ’Name’, ’Mixed-Language’, ’Other’, and ’Location’. The primary goal of the dataset is to improve techniques for language identification and classification, particularly for Kannada-English code-mixed texts.
@@
    Tamil Dataset: The Tamil dataset comprises 17,568 tokens, created using a methodology similar to that employed for the Kannada and Tulu datasets. It is divided into six categories and is designed to facilitate a range of NLP tasks tailored to the Tamil language.
@@
    Malayalam Dataset: This dataset consists of 25,035 tokens classified into 7 categories: ’Number’, ’Mixed’, ’English’, ’Location’, ’Name’, ’sym’ (for sentence boundaries), and ’Malayalam’. This dataset offers extensive coverage for NLP tasks and includes the ’Number’ category for numerical values, akin to the structure of the other provided datasets.
We participated in shared tasks based on two languages, namely, Kannada and Tamil. The test dataset size for Kannada is 2502 words. The test dataset size for Tamil is 2024 words.
@@ 5.1 Why Prompting?
Prompting [38] to solve a word-level classification problem often arises from the need to accurately identify and categorize individual words within texts that exhibit code-mixing or multilingual content. Next we discuss the reasons why the problem of language identification is tried via prompting through GPT-3.5 Turbo:
@@
  • Accuracy in Language Processing: For effective natural language processing (NLP), identifying the language of each word enhances the accuracy of subsequent tasks [40]. It ensures that language-specific models or algorithms are applied correctly.
@@ 5.2 Prompt Engineering-Based Approach
We used the GPT-3.5 Turbo model via prompting [^1] https://platform.openai.com/docs/models/gpt-3-5-turbo to solve the classification task in Zero-shot mode. After the prompt is provided to the LLM, the following steps occur internally while generating the output. GPT-3.5 Turbo follows a decoder-only architecture. So based on [3, 26, 43], we list these steps, summarizing the prompting approach using GPT-3.5 Turbo. The set of steps [26, 43] for GPT-3.5 Turbo [3] is as follows:
@@
Corresponding to the two distinct prompts (for Kannada and Tamil) the two distinct figures are stated (Figure 2 and Figure 3).
@@ 6 Results
Table 1: Comparison of various metrics for Word level identification in code-mixed languages in two languages-Tamil, Kannada. The team name is TextTitans and the username is roydanik18. All metric values have been reported in this table.
Metric Tamil Kannada
Macro F1 0.3312 0.4493
Macro Precision 0.3259 0.5474
Macro Recall 0.3657 0.4241
Weighted F1 0.7022 0.6725
Weighted Precision 0.7559 0.7191
Weighted Recall 0.6689 0.6994
Accuracy 0.6689 0.6994
Table 1 presents metrics comparing the performance of two language identification models, one for Tamil and the other for Kannada. Our Team ranked in the 10th position for both tasks. Here’s a detailed discussion of each metric.
Next, we discuss the metric values as well as their corresponding analysis to explain the results in the Table  1. For Tamil language, the macro F1 score is 0.3312. This suggests that the model achieves a balanced performance in terms of precision and recall for Tamil language identification. However, it indicates there is room for improvement in correctly identifying both positive and negative instances. For Kannada, the macro F1 score is 0.4493. This score is higher compared to Tamil, indicating a better overall balance between precision and recall for Kannada language identification. The model for Kannada performs better in correctly classifying instances across the dataset.
@@
The weighted precision, recall, and f1-scores being higher than the macro precision, recall, and f1-scores shows that dataset likely has an imbalance, with some classes having many more samples than others. The weighted F1 score takes this into account by giving more importance to the performance on larger classes. The model is performing well on the classes that contribute the most to the overall accuracy. This could mean that it is effectively identifying the majority classes but may struggle with minority classes for both languages.
Weighted precision being higher than weighted recall suggests that the model performs better on the more frequent classes in the dataset. This means it is more effective at correctly identifying positive instances for these majority classes for both datasets. For Kannada dataset, a higher macro precision than recall may suggest that the model is conservative in its positive predictions, prioritizing accuracy over completeness. For Tamil dataset, a higher macro recall than precision suggests that while the model is effective at capturing relevant instances, it may not be very reliable in its predictions.
@@ 7 Conclusion
Moving forward, further refinements in model training and dataset augmentation could enhance the performance of language identification systems for both Tamil and Kannada. Future research efforts may focus on incorporating additional linguistic features, optimizing model architectures, and expanding datasets to include more diverse linguistic variations and challenges. In conclusion, this study underscores the importance of tailored approaches in language identification, particularly in multilingual settings like India where linguistic diversity is prominent. By advancing the capabilities of language identification models through innovative methodologies such as GPT-3.5 Turbo via prompting, we contribute to the broader goal of improving language processing technologies for diverse and under-resourced languages, fostering more accurate and inclusive natural language understanding systems. Future work would focus on improving the prompts to improve accuracy on the language identification task.